a-8/ 12 = 2a +5/3
The whole part 8-12 is divided by 12 and the whole part 2a +5 is divided by 3. Dont you cross multiply to get
a-8 = 8a + 20 .4
then what?

Answer by Alan3354(69443)

a-8/ 12 = 2a +5/3
The whole part 8-12 is divided by 12 and the whole part 2a +5 is divided by 3. Dont you cross multiply to get
a-8 = 8a + 20 .4
then what?
===============
( a-8)/12 = (2a +5)/3
3(a-8) = 12(2a+5)
3a-24 = 24a+60
-84 = 21a
a = -4